On average, indoor cats live to 15 years old with a standard deviation of 2.3 years. Suppose that the distribution is normal. Let X = the age at death of a randomly selected indoor cat. Round answers to 4 decimal places where possible.

a. What is the distribution of X? X ~ N(
,
)

b. Find the probability that an indoor cat dies when it is between 13.9 and 16.1 years old.


c. The middle 20% of indoor cats' age of death lies between what two numbers?
Low:
years
High:
years
1
Expert's answer
2019-11-11T11:12:55-0500

a) Given that "\mu=15 \ years, \sigma=2.3\ years"

"X\sim N(\mu, \sigma^2)"

"X\sim(15, 2.3^2)"

b) "X\sim N(\mu, \sigma^2)." Then


"Z={X-\mu \over \sigma}\sim N(0, 1)"

"\mu=15 \ years, \sigma=2.3\ years"


"P(13.9<X<16.1)=P(X<16.1)-P(X<13.9)="

"=P(Z<{16.1-15 \over 2.3})-P(Z<{13.9-15 \over 2.3})\approx"

"\approx0.683768-0.316232\approx0.3675"

"P(13.9<X<16.1)=0.3675"

c)


"P(\mu-\delta<X<\mu+\delta)=0.2"

"P(Z<{\mu+\delta-\mu \over \sigma})-P(Z<{\mu-\delta-\mu \over \sigma})=0.2"

"P(Z<{\delta \over \sigma})-P(Z<-{\delta \over \sigma})=0.2"

"P(Z<-{\delta \over \sigma})={1-0.2 \over 2}=0.4"

"-{\delta \over \sigma}\approx-0.253348"

"\delta\approx2.3\cdot0.253348=0.5827004"

"\mu-\delta\approx15-0.5827004\approx14.4173"

"\mu+\delta\approx15+0.5827004\approx15.5827"

"Low:14.4173 \ years"

"High:15.5827\ years"